How many years will it take for $14,000 to grow to $31,000 if you can earn an average rate of 7.8% per year?

NOTE: Please enter the answer rounded to the nearest one digit after the decimal al point.
Answer #1

We use the formula:  
A=P(1+r/100)^n
where   
A=future value
P=present value  
r=rate of interest
n=time period

31000=14000*(1.078)^n

(31000/14000)=(1.078)^n

Taking log on both sides;

log (31000/14000)=n*log 1.078

n=log (31000/14000)/log 1.078

=10.6 years(Approx)